OpenStudy (anonymous):

Find the volume of a cone with a diameter of 9 inches and a height that is 5 times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (5 points) Select one: a. 47.1 in3 b. 78.51in3 c. 476.89 in3 d. 1430.66 in3

OpenStudy (anonymous):

Just use the formula V = 1/3(πr²)h.

OpenStudy (anonymous):

Since diameter is radius doubled, d = 2r Since the diameter is 9, 9 = 2r 9 ÷ 2 = 2r ÷ 2 4.5 = r In this case, height is 5 times the radius, so h = 5r r = 4.5 h = 5(4.5) h = 22.5 Now that you have your values, just plug them into the formula that I gave you. You can figure out the rest. :)